Nungesser Corporation has issued bonds that have a 9 percent coupon rate, payable semiannually. The bonds mature in 6 years, have a face value of $1,000, and a yield to maturity of 8.5 percent. Current Market Price = PV(Rate,Nper,PMT,FV) 1). What is the price of the bonds? $1,023.13 2). What is the current yield? 8.80% 3). What is the capital gains yield? -0.30% 4). These bonds sell at a. par b. a premium c. a discount

Please type out all responses as I am on mobile no grid style or excel please
Answer #1

Coupon Payment = 9% * 1,000 = 90

Semi Annual Payment = 90/ 2 = 45

Time to maturity = 6 years

Semi annual periods to maturity = 6 * 2 = 12

FV = 1,000

YTM = 8.5%

Semi Annual yield = 8.5%/ 2 = 4.25%

PV = PMT * (1 - (1 + Interest rate)-time)/ Interest rate + Maturity Price/ (1 + interest rate)time

PV = 45 * (1 - (1 + 4.25%)-12)/ 4.25% + 1,000/ (1 + 4.25%)12

PV = 45 * (1 - 0.606858)/ 4.25% + 1,000/ 1.65

PV = 416.27 + 606.86

PV = 1,023.13

Part b

Current yield = Coupon payment/ Price

Current yield = 90/ 1,023.13

Current yield = 8.80%

Part C

Capital gain yield = YTM - Current yield

Capital gain yield = 8.50% - 8.80%

Capital gain yield = -0.30%

Part 4

As coupon rate is more than YTM, the bond is a premium bond